# Blockchain Address Poisoning

### v1

@@ 5.1 Detection results
We present the summary statistics for poisoning transfers in Table 1 and confirmed payoff transfers in Table 2
@@
Total Loss Min Median Avg. Max Std. dev Count Victims (USD) (USD) (USD) (USD) (USD) (USD) Ethereum 79,344,412 0.001 2,160 45,653 20,000,000 521,663 1,738 1,502 BSC 4,490,804 0.000 85 1,164 279,489 8,366 4,895 4,004
@@
Attackers on Ethereum receive 1,738 transfers from 1,502 victims, totaling nearly 80M USD (avg. 45,853 USD, median 2,169 USD, with high variance), with a success rate is 0.01%8880.13% spam click-through rate in Twitter [9]. Some victims sent assets to the attacker’s address(es) multiple times. For instance, one victim sent 1.999M and 2M USDC within 10 blocks without noticing they were being phished.
@@ 5.2 Detection evaluation
Manual evaluation on successful cases. Statistically, the probability of getting a false positive is exceedingly low because 1) we only look at transfers where the victim has interacted with two similar addresses, and 2) we observe poisoning transfers between an intended and a payoff transfer take place. Nevertheless, two of the authors also manually verified transfers labeled as successful payoff transfers. We sorted payoff transfers by decreasing the amount and checked the 30 largest transfers for both chains. We visit each victim’s page on external blockchain scanning websites (e.g., Etherscan and BSCscan) and manually confirmed 1) the intended (original) transfer, 2) the poisoning transfer, and 3) the payoff transfer. We verified that all 30 cases are indeed payoff transfers for both Ethereum and BSC (i.e., no false positives). Those 30 cases account for 53.5M (67%) of Ethereum losses and 2M (45%) of BSC losses. For accountability purposes, we list the top 30 successful cases for Ethereum and BSC in Appendix G.
Comparison with previous work. We directly compare our detection results by focusing on 12 and 16 months of data to align with Ye et al.’s work [46] and Guan and Li’s [10] work, respectively. We detect more payoff transfers than Ye et al. We find 1,059 successful cases with 36.6M USD in losses while Ye et al. found 560 successful cases with at most 28M USD in losses, including two separate (non-address poisoning) attacks they investigate. In terms of poisoning transfers, we identify 4.9M zero-value token transfers, compared to 21M for Ye et al. It is unclear whether the difference comes from 1) false positives on Ye et al.’s side, (i.e., benign zero-value token transfers), or 2) false negatives on our side.
We find a similar total loss on Ethereum (during the same period) as Guan and Li [10], but we detect more than 2.2M counterfeit token transfers, mostly from tokens with modified symbols (refer to a more detailed comparison in Appendix A).
@@
Window range. To make the chain scanning process efficient, we assume that the attack happens in the next 20 minutes (100 blocks in Ethereum) after the original transfers. We evaluate this assumption by collecting data over 200,000 blocks starting from 16,950,603 (Apr. 1st, 2023), using both a 100-block range and a 200-block range, and comparing results. Both ranges identify the same number of successful cases (90). The 200-block range only identifies two additional tiny transfers besides the 250 tiny transfers identified with a 100-block range. With a 100-block range, we find 102,154 zero-value transfers; extending to 200 blocks only identifies an additional 27 zero transfers. Finally, a 100-block window captures 704,998 counterfeit token transfers, whereas a 200-block window brings this up to 706,332. In short, doubling the window size (20 minutes to 40 minutes) does not dramatically increase the number of attacks we discover. We believe that collecting other transfers within the same poisoning transaction (§4.1: Step 2) has significantly helped the coverage.
@@ 5.3 Clustering results
Before clustering, we attempt to remove addresses (bots) that copy transactions by setting an attack ratio threshold, as discussed in §4.2. We choose a threshold of 0.5, which 1) avoids erroneous merge of large clusters and 2) only removes 1.49% of all transfers. We provide a detailed explanation of this threshold selection and demonstrate the robustness of our clustering results (through temporal clustering) in Appendix D.
Table 3 describes the number of attack instances and assigns Group IDs sorted by the number of lookalike addresses in each group on Ethereum. The notation follows §3. We only list the top eight biggest groups because the size of the eighth group is significantly smaller than the top seven (i.e., less than 100,000 lookalike addresses with a few successful cases). Different groups exhibit different attack strategies; not all groups use three poisoning transfers. Only Group 1 uses tiny transfers. Groups 3, 4, and 7 do not have any counterfeit token contracts, $CT$, and thus do not use any counterfeit-token transfers, and instead focus on zero-value transfers.
We additionally collect contract bytecode from our Ethereum node and list the number of distinct contracts in a parenthesis. In Group 2, for example, all $CT$ and $AC$ are identical, indicating the reuse of the same source code (but with different addresses). Given that the contract bytecode changes even with a slight alteration in the source code, the high number of identical contracts corroborates the validity of our clustering techniques (especially since our clustering does not rely on those contracts).
Table 3: Attack group general statistics. Notations follow Fig. 2). The distinct number of contracts is in parentheses.
@@
We also investigate the number of poisoning transfers over time. Rather than focusing on the aggregate number, we present the number of attacks from each group. Figure 4 illustrates the number of poisoning transfers (not transactions) per week from the top seven groups. We started to observe attacks from Dec. 2022, starting with Group 7. Early in 2023, Groups 3 and 4 enter the market and Group 1 comes around Feb. 2023. As we will see in Table 4, we could not confirm an early mover advantage (i.e., later comers are still profitable). The variation in the attack strategies and the active periods across attack groups further supports our clustering results.
@@ 6.1 Targeted population
We first look at victim profiles, so that we can better understand who the attackers target, and eventually propose defenses. To make the attack more profitable, we conjecture that attackers are more likely to target rich addresses. As different attack groups may employ different strategies, we look at victims on a group-by-group basis on Ethereum. For this purpose, we maintain an Ethereum archive node that retains all historical account states. Specifically, we randomly sample 1,000 victims from each of the top seven attack groups. For each victim, we retrieve the sum of their three stablecoins (USDT, USDC, DAI) balance just before when the victim receives the first poisoning transfer from each attack group. We also create a baseline group by randomly sampling 1,000 USDT users and looking up their balance before the most recent transfer. Figure 5 illustrates the cumulative distribution function (CDF) of the total stablecoin balance for the victims in each of the seven attack groups and the baseline USDT users. All attack groups tend to target addresses that have significantly more stablecoins than regular USDT users. Groups 6 particularly target rich accounts given that most of the victims own more than 10,000 USD at the time of the attack. Targeting rich accounts seems successful since we detect 83 payoff transfers with more than 100,000 USD.
@@ 6.2 Attack strategies
First, we discover that the attacker tends to imitate addresses ($\mathcal{R}$) from centralized entities. By manually looking at the list of addresses in $\mathcal{R}$ on Ethereum that the attacker imitates to generate $L$ the most, we find that those addresses often belong to centralized exchanges (e.g., Coinbase, Binance, Bybit) probably because they are one of the most active accounts. However, most addresses are hot wallets, with which normal users typically do not interact. Indeed, we do not find any successful attack involving those hot wallets. This appears to be a waste of time and money, as excluding such addresses would reduce address generation and transaction costs. We have more detailed information in Appendix E.
@@
Figure 7: Number of lookalike addresses for which the first $x$- and the last $y$-digit match the intended recipient address.
@@
Figure 8: Distributions of the lookalike address similarity (for the top 4 attack groups): y-axis is on a log scale
@@ 6.3 When is the attack successful?
We next look into which conditions are satisfied for an attack to be successful. We test our hypothesis that the more similar the lookalike address is to the intended recipient address, the more likely the victim will make a mistake. In Figure 9, the bubble size represents the success ratio (the number of successes to the number of lookalike addresses) depending on similarity. We do not consider the number of attack attempts for each victim. While $(x,y)=(7,10)$ and $(x,y)=(9,8)$ have high success rate, there is no strong linear correlation between similarity and success ratio in general. This result suggests that the attack may have been more profitable to generate a larger number of slightly less convincing lookalike addresses and conduct more attacks.
Figure 9: Success probability and address similarity. The first $x$- and the last $y$-digit match the intended recipient address.
We next focus on successful cases and examine how address similarity (or attack timing) relates to success. Figure 10 (left) is the number of lookalike addresses targeting each victim. In 79% of the successful attacks on Ethereum, multiple lookalike addresses target one victim, indicating competition among attackers. This competition allows us to investigate how a victim selects one particular lookalike address (i.e., a winner $L_{\text{win}}$) over a pool of other lookalike addresses (i.e., victims’ perception). We focus on two aspects: how similar the winner $L_{\text{win}}$ is to $R$; and how early the winning address appears, compared to other lookalike addresses. For each successful attack, we extract the list of corresponding lookalike addresses $\mathcal{L}$, then check how $L_{\text{win}}$—the recipient of the payoff transfer—ranks in these two metrics. A smaller rank implies a higher similarity; and an earlier attack, respectively. Figure 10 shows that, while the number of competitors (left) decreases somewhat linearly, a disproportionately large proportion of the most similar addresses (center), and/or early-comers (right) tend to win the competition. In other words, in times of stronger competition, being the first and/or having a lookalike address most similar to the intended recipient seems to yield measurable benefits.
While those results provide insights from chain activities, the success also relies on what wallets or chain scanning services victims use. Appendix 8 compares how poisoning attacks appear in those services’ UI and qualitatively discusses when and where victims are most likely to fall for the attack.
@@ 6.4 Benefit-cost analysis for attack groups
In general, the top seven groups are largely profitable. Despite the high variance in payoff amounts as shown in Table 2, those groups manage to amortize the revenue across many payoff transfers while some small groups do not. Given the attack scale (i.e., Group 1 has invested over 8 million USD), some groups appear to be well-funded cyber-crime entities. However, none of the attack instances comes from any sanctioned entities in the SDN list [41].
@@ 7.1 Mathematical formulation
We assume that the attacker wants to match the first $a$ and the last $b$ digits of the intended recipient and maximize the total number of digits $a+b=d$, with no preference for $a$ over $b$. For instance, the attacker is indifferent between $(a,b)=(3,6)$ or $(4,5)$ as long as $(a\geq 3,b\geq 4)$ holds, per §4.1. The attacker has the set of target addresses ($\mathcal{R}$) they want to imitate.
@@
$p=1-\left(\frac{16^{d}-1}{16^{d}}\right)^{r}\ ,$ (1)
which reduces to $p=1/16^{d}=2^{-4d}$ when $r=1$ (i.e., a single address is targeted).
The probability of finding a collision after $k$ trials follows the geometric distribution $P(X=k)=p\cdot(1-p)^{k-1}$, and the expected number of trials for a collision is $E[X]=1/p$. Using a first-order Taylor approximation in Eqn. (1) yields $p\approx r/16^{d}$, and the resulting expected number of trials is $E[X]\approx 2^{4d}/r$. In short, the collision probability linearly increases with the size of the target address space $r$, and, accordingly, the expected number of trials linearly decreases with $r$.
@@ 7.2 Simulating address generation
We implement 1) a naive version using Python with the web3py [^10] https://github.com/ethereum/web3.py library; and 2) an optimized version leveraging CPU multiprocessing using Python with the multiprocessing, coincurve, [^11] https://github.com/ofek/coincurve, Python bindings for a heavily optimized C library for elliptic curve operations. and pycryptodome [^12] https://github.com/Legrandin/pycryptodome/, the default hashing package used by web3py. libraries. In addition, we use two popular vanity address-generating software: Vanity-ETH, [^13] https://vanity-eth.tk/ a Javascript-based web application, and Profanity2, [^14] https://github.com/1inch/profanity2 a C++-based GPU implementation. We elect, for simplicity, not to check whether the generated address is in the set of target addresses $\mathcal{R}$, which will produce an optimistic bound on performance, and thus a conservative bound on the attackers’ implied hardware capabilities.
We test our implementations on two hardware configurations, a Macbook Air (OS Version 13.1) with an 8-core Apple M1 base chip; and a server workstation (Ubuntu 22.04 LTS) with an Intel(R) Xeon(R) Silver 4214 CPU and an Nvidia RTX 8000 GPU. We run our simulations on the CPU on both machines and on the GPU on the second machine. We summarize our simulation results in Table 5.
The naive Python implementation which models what might be an expected “first-attempt” by an attacker produces in the order of $10^{3}$$10^{4}$ addresses per second. The more optimized implementation produces an order of magnitude improvement ($10^{4}$$10^{5}$). The Vanity-ETH implementation is much slower than the standalone clients. The CUDA implementation from Profanity2 is three orders of magnitude faster than our best CPU implementation with roughly half a billion APS. The Profanity2 developers self-report about a billion APS on similarly priced consumer hardware. However, the overhead imposed by checking set inclusion in $\mathcal{R}$ should be much higher in the GPU implementation.
@@
Based on the distribution of the matching digits $d$ for each attack group, we next attempt to infer the attacker’s hardware capability. We define two units: CPU-day and GPU-day, the number of addresses the attacker can generate in one day per machine with a CPU and a GPU, respectively. From the above benchmarks, 1 CPU-day is $3.98\times 10^{10}$ addresses and and 1 GPU-day is $4.46\times 10^{13}$, respectively. In the process of producing a $d$-digit match, the attacker can generate many addresses matching $d^{\prime}<d$ digits in $\mathcal{R}$. Therefore, estimating the amount of computation needed to generate the maximum match $d$ for each group suffices.
As shown in Figure 8, Group 1 produces a maximum 20-digit match while other groups (Groups 2, 3, and 4) achieve 14-digit matches. We calculate the number of expected trials for $d=20$ and, based on Table 3, $r=10^{6}$. We estimate that Group 1 have used $3.0\times 10^{7}$ CPU-days or 27,093 GPU-days.
Group 1 would have needed to run over 41,000 CPUs around the clock over our whole two years of measurements to generate their lookalike addresses. We thus suspect they perform a substantial amount of their computation via GPUs. Group 1 could also have developed specialized hardware for address generation (e.g., ASICs), which would further reduce generation time. Other groups ($d=14$) only used 1.81 CPU-days, or $1.6\times 10^{-3}$ GPU-days.
To get a rough idea of the economic cost of computation, we use the AWS pricing model for similar GPUs and CPUs [^15] GPUs: https://aws.amazon.com/ec2/instance-types/g4/ and CPUs: https://aws.amazon.com/ec2/pricing/on-demand/. For GPU, an hourly rate (yearly plan) for g4dn.16xlarge (NVIDIA T4 tensor) is \$2.612 per hour; \$62.69 for one GPU-day. For CPU, an hourly rate (on-demand) for r6i.4xlarge is \$1.008; \$24.19 for one CPU-day. Using these numbers, we estimate that Group 1 would have paid (at most) 1.7M USD to generate a 20-digit match, which is well below its 4M USD profit described in Table 4. Furthermore, these numbers are likely higher than the costs incurred by self-hosting hardware. For other groups working with $d=14$, CPU costs ($1.81\times 24.19\approx 43$ USD) are dominated by transaction costs.

### v3

@@ 5.1 Detection results
We present the summary statistics for poisoning transfers in Table 1 and confirmed payoff transfers in Table 2.
@@
Total Loss Min Median Avg. Max Std. dev Count Victims (USD) (USD) (USD) (USD) (USD) (USD) Ethereum 79,344,412 0.001 2,160 45,653 20,000,000 514,142 1,738 1,502 BSC 4,490,804 0.000 85 1,164 279,489 7681 4,895 4,004
@@
Attackers on Ethereum receive 1,738 transfers from 1,502 victims, totaling nearly 80M USD (avg. 45,853 USD, median 2,169 USD, with high variance), with a success rate is 0.01%9990.13% spam click-through rate in Twitter [9]. Some victims sent assets to the attacker’s address(es) multiple times. For instance, one victim sent 1.999M and 2M USDC within 10 blocks without noticing they were being phished.
@@ 5.2 Detection evaluation
Evaluation with external dataset. For our detector evaluation, we obtained from Guan and Li [10] their evaluation dataset and detection results. [^10] The authors kindly made the evaluation open source recently: https://github.com/DS2L/Poison-Hunter. The data consists of 5,890 lookalike addresses and 1,154 benign (non-lookalike) addresses. The authors include lookalike addresses only when two sources (Forta and Etherscan) report them and label DeFi entities as benign based on the Etherscan Label Cloud. [^11] https://etherscan.io/labelcloud We only use our results from Nov. 2022 to Feb. 2024 to align with their experiment. We discovered and manually confirmed that 45 addresses are mislabeled as lookalike—there are no corresponding poisoning transfers. We have contacted Guan and Li, who agreed to remove these mislabeled entries from their dataset. After removing mislabeled entries, our detector achieves 97.7% accuracy, 100% precision, 97.2% recall, and a 98.58% F1-score. In absolute terms, we identify 5,681 lookalike addresses with 164 false negatives and 0 false positives.
We investigate those 164 false negatives using Etherscan. A full discussion of the whole analysis can be found in Appendix A. Here we summarize the high-level results. In 87 instances of false negatives, we fail to identify the lookalike address because the victim engaged in a benign transfer that resembles a poisoning transfer. For instance, the victim may have sent a legitimate zero-value transfer to the intended address for testing. Our detector cannot distinguish between that case and that of an attacker creating a zero-value transfer appearing to originate from the victim, and sent to a lookalike address. In such a case, our detector conservatively does not classify the event as an attack, and tags the corresponding lookalike address as legitimate. 23 false negatives are due to attacks being outside our time-window range (i.e., attackers target the victim after 20 minutes). 16 false negatives come from attackers targeting other tokens (e.g., TrueUSD, BUSD) while we only look at major stablecoin transfers. The rest (38 false negatives) can be primarily attributed to data collection quirks. Importantly, our choice of similarity threshold $(3,4)$ or that our detector only considering senders as potential victims is not the cause of any of the false negatives.
Guan and Li’s detector discovers 144 new addresses we did not catch, while our detector captures 96 addresses that Guan and Li’s did not. Guan and Li’s detector employs additional checks using attack characteristics (e.g., checking a transaction initiator) and filters out benign transfers that resemble poisoning transfers (i.e., mitigating the first case of false negatives we identified above). However, Guan and Li’s approach (slightly overfitting to known attacks) may miss poisoning transfers when the attacker modifies their attack strategy (e.g., using a lookalike address to launch a poisoning transaction).
@@
Manual evaluation on successful cases and attack contracts. Statistically, the probability of getting a false positive is exceedingly low because 1) we only look at transfers where the victim has interacted with two similar addresses, and 2) we observe poisoning transfers between an intended and a payoff transfer take place. Nevertheless, two of the authors also manually verified transfers labeled as successful payoff transfers. We sorted payoff transfers by decreasing the amount and checked the 30 largest transfers for both chains. We visited each victim’s page on external blockchain scanning websites, Etherscan and BSCscan, and manually confirmed 1) the intended (original) transfer, 2) the poisoning transfer, and 3) the payoff transfer and their order. We verified that all 30 cases are indeed payoff transfers for both Ethereum and BSC (i.e., no false positives). Those 30 cases account for 53.5M (67%) of Ethereum losses and 2M (45%) of BSC losses.
We perform the same operation for the contracts used for poisoning transfers ($AC$). We sort $AC$ based on the number of poisoning transfers, manually look at the top 30 $AC$ for each chain on external blockchain scanning websites, and confirm at least one poisoning transfer (and its corresponding intended transfer on the victim’s page). We verify that all the contracts are indeed attack contracts, and they account for 4.9M (28%) on Ethereum and 210.1M (83%) on BSC.
@@
Parameter selection. We set key assumptions about the window size (100 blocks) and the similarity threshold (3, 4) to make the chain scanning process efficient and reduce false positives. We evaluate those assumptions by collecting data over 200,000 blocks starting from 16,950,603 (Apr. 1st, 2023), and performing a detection with a window size of 200 and similarity score of (3, 3). Table 3 compares the different parameters and summarizes the results on the number of poisoning transfers, lookalike addresses, and payoff transfers.
The 200-block range captures 1,363 (0.17%) more poisoning transfers, 1,139 (0.27%) more lookalike addresses, and the same number of payoff transfers. Hence, doubling the window size does not dramatically increase the number of attacks we discover. We believe that collecting other transfers within the same poisoning transaction (§4.1: Step 2) has significantly helped the coverage.
When we select the threshold of (3, 3) instead of (3, 4), we capture 2,937 (0.36%) additional poisoning transfers and 972 more lookalike addresses (0.23%). However, if we decrease the threshold by 1, we are $\approx\sqrt{16}=4$ times more likely to find benign lookalike addresses. One additional payoff transfer identified with a threshold of (3, 3) was indeed a false positive. In short, changing those parameters does not appear to outweigh the cost of data collection and false positives.
Table 3: Detection results with different parameters. T: tiny, Z: zero-value, C: counterfeit token transfer, L: lookalike address
Window size Similarity Poisoning (T, Z, C) L Payoff
100 (3, 4) 807,402 (250, 102,154, 704,998) 427,317 90
200 (3, 3) 808,765 (252, 102,181, 706,332) 428,456 90
100 (3, 3) 810,339 (265, 102,363, 707,711) 428,289 91
@@ 5.3 Clustering results
Before clustering, we attempt to remove addresses (bots) that copy transactions by setting an attack ratio threshold, as discussed in §4.2. We choose a threshold of 0.5, which 1) avoids erroneous merge of large clusters and 2) only removes 1.49% of all transfers. We provide a detailed explanation of this threshold selection in Appendix E.
In total, we identify 49 groups (with more than one lookalike address), corresponding to 97.4% of poisoning transactions, corroborating our detection results. Table 4 describes the number of attack instances and assigns Group IDs sorted by the number of lookalike addresses in each group on Ethereum, following the notation in §3. We only list the top eight groups because the eighth group is significantly smaller than the top seven (i.e., less than 100,000 lookalike addresses with a few successful cases). Different groups exhibit different attack strategies; not all groups use all three poisoning types. Only Group 1 uses tiny transfers. Groups 3, 4, and 7 do not use any counterfeit token transfers, and instead focus on zero-value transfers. Additionally, we collect contract bytecode from our Ethereum node and list the number of distinct contracts in parentheses. In Group 2, for example, all $CT$ and $AC$ are identical, indicating the same source code reuse (but with different addresses). Given that the contract bytecode changes even with a slight alteration in the source code, the high number of identical contracts corroborates the validity of our clustering techniques (especially since our clustering does not rely on those contracts).
Table 4: Attack group statistics. Notations follow Fig. 2. The distinct number of contracts is in parentheses.
@@
We also investigate the number of poisoning transfers over time. Rather than focusing on the aggregate number, we present the number of attacks from each group. Figure 4 illustrates the number of poisoning transfers (not transactions) per week from the top seven groups. We started to observe attacks from Dec. 2022, starting with Group 7. Early in 2023, Groups 3 and 4 enter the market and Group 1 comes around Feb. 2023. While several top groups stop operations around May 2024 (in Figure 4, Ethereum), the attack is still active today; we observe a drastic increase in BSC in June 2024 (in §5.1). As we will see in Table 5, we could not confirm an early mover advantage (i.e., later comers are still profitable). The variation in the attack strategies and the active periods across attack groups further supports our clustering results.
All in all, we consistently see behavioral heterogeneity across the groups we identified, such as distinct attack strategies, identical contract reuse, and different activity periods. In Appendix E, we also demonstrate the robustness of our clustering results through temporal clustering; most attack groups stay stable/distinct over time.
@@ 6.1 Targeted population
Attackers appear to select their targets in advance because the attack typically happens right after the target’s original transfer (mostly within 20 minutes, as shown in §5.2). We further investigate whether attackers disproportionately target certain populations compared to the average stablecoin user.
To make the attack more profitable, we conjecture that attackers are more likely to target rich addresses. As different attack groups may employ different strategies, we look at victims on a group-by-group basis on Ethereum. For this purpose, we maintain an Ethereum archive node that retains all historical account states. Specifically, we randomly sample 1,000 victims from each of the top seven attack groups. For each victim, we retrieve the sum of their three stablecoins (USDT, USDC, DAI) balance just before when the victim receives the first poisoning transfer from each attack group. We also create a baseline group by randomly sampling 1,000 USDT users and looking up their balance (for three stablecoins) before the most recent transfer. Figure 5 illustrates the cumulative distribution function (CDF) of the total stablecoin balance for the victims in each of the seven attack groups and the baseline USDT users. All attack groups tend to target addresses that have significantly more stablecoins than regular USDT users. Groups 6 particularly target rich accounts given that most of the victims own more than 10,000 USD at the time of the attack. Targeting rich accounts seems successful since we detect 83 payoff transfers with more than 100,000 USD.
@@ 6.2 Attack strategies
First, we discover that the attacker tends to imitate addresses ($\mathcal{R}$) from centralized entities. By manually looking at the list of addresses in $\mathcal{R}$ on Ethereum that the attacker imitates the most, we find that those addresses often belong to centralized exchanges (e.g., Coinbase, Binance, Bybit), probably because they are one of the most active accounts. However, most addresses are hot wallets, with which normal users typically do not interact. Indeed, we do not find any successful attack involving those hot wallets. This appears to be a waste of time and money, as excluding such addresses would reduce address generation and transaction costs. We have more detailed information in Appendix F.
@@
Figure 7: Number of lookalike addresses for which the first $x$- and the last $y$-digit match the intended address.
@@
Figure 8: Distributions of the lookalike address similarity (for the top 4 attack groups): y-axis is on a log scale.
@@ 6.3 When is the attack successful?
We next look into which conditions are satisfied for an attack to be successful. We test our hypothesis that the more similar the lookalike address is to the intended address, the more likely the victim will make a mistake. In Figure 9, the bubble size represents the success ratio (the number of successes to the number of lookalike addresses) depending on similarity. We do not consider the number of attack attempts for each victim. While $(x,y)=(7,10)$ and $(x,y)=(9,8)$ have high success rate, there is no strong linear correlation between similarity and success ratio in general. This result suggests that the attack may have been more profitable to generate a larger number of slightly less convincing lookalike addresses and conduct more attacks.
Figure 9: Success probability and address similarity. The first $x$- and the last $y$-digit match the intended address.
We next focus on successful cases and examine how address similarity or attack timing relates to success. Figure 10 (left) shows the number of lookalike addresses targeting each victim. In 79% of the successful attacks on Ethereum, multiple lookalike addresses target one victim, indicating competition among attackers. This competition allows us to investigate how each victim selects one particular lookalike address, a winner $L_{\text{win}}$, over a pool of other lookalike addresses, while partially controlling for individual user attributes. We focus on two aspects: how similar the winner $L_{\text{win}}$ is to $R$; and how early the winning address appears, compared to other lookalike addresses. For each successful attack, we extract the list of corresponding lookalike addresses $\mathcal{L}$, then check how $L_{\text{win}}$—the recipient of the payoff transfer—ranks in these two metrics. A smaller rank implies a higher similarity and an earlier attack, respectively. Figure 10 shows that, while the number of competitors (left) decreases somewhat linearly, a disproportionately large proportion of the most similar addresses (center), and/or early-comers (right) tend to win the competition. In other words, in times of stronger competition, being the first and/or having a lookalike address most similar to the intended recipient seems to yield measurable benefits.
While those results provide insights from chain activities, the success also relies on what wallets or chain scanning services victims use. Appendix G compares how poisoning attacks appear in those services’ UI and qualitatively discusses when and where victims are most likely to fall for the attack.
@@ 6.4 Benefit-cost analysis for attack groups
In general, the top seven groups are largely profitable. Despite the high variance in payoff amounts as shown in Table 2, those groups manage to amortize the revenue across many payoff transfers while some small groups do not. Given the attack scale (i.e., Group 1 has invested over 8 million USD), some groups appear to be well-funded cybercrime entities. However, none of the attack instances comes from any sanctioned entities in the SDN list [42].
@@ 7.1 Mathematical formulation
We assume that the attacker wants to match the first $a$ and the last $b$ digits of the intended address and maximize $a+b=d$, with no preference for $a$ over $b$. For instance, the attacker is indifferent between $(a,b)=(3,6)$ or $(4,5)$ as long as $(a\geq 3,b\geq 4)$ holds, per §4.1. The attacker has a set of target addresses ($\mathcal{R}$) they want to imitate as discussed in §6.1.
@@
$p=1-\left(\frac{16^{d}-1}{16^{d}}\right)^{r}\ .$ (1)
Using a first-order Taylor approximation yields $p\approx r/16^{d}$.
The probability of finding a collision after $k$ trials follows the geometric distribution $P(X=k)=p\cdot(1-p)^{k-1}$, and the expected number of trials for a collision is $E[X]=1/p\approx 2^{4d}/r$. In short, the expected number of trials exponentially increases with the the number of matched digits, and linearly decreases with the size of the target address space.
@@ 7.2 Simulating address generation
For address generation software, we have tested several combinations of libraries: web3py [^13] https://github.com/ethereum/web3.py., coincurve [^14] https://ofek.dev/coincurve/, a set of Python bindings for a heavily optimized C library for elliptic curve operations., pycryptodome [^15] https://github.com/Legrandin/pycryptodome/, the default hashing package used by web3py., with/without concurrency, and multiprocessing. We also implement the address generation scripts in Go (with concurrency), which yields results similar to those obtained with Python. We additionally use two popular vanity address software: Vanity-ETH, [^16] https://vanity-eth.tk/ a Javascript-based web application, and Profanity2, [^17] https://github.com/1inch/profanity2 a C++-based GPU implementation. We elect, for simplicity, not to check whether the generated address is in the set of target addresses $\mathcal{R}$, which will produce an optimistic bound on performance, and thus a conservative bound on the attackers’ implied hardware capabilities.
We test our implementations on two hardware configurations, a MacBook Air (OS Version 13.1) with an 8-core Apple M1 base chip, and a server workstation (Ubuntu 22.04 LTS) with an Intel(R) Xeon(R) Silver 4214 CPU and an Nvidia RTX 8000 GPU. Other hardware settings (e.g., different generations of Mac chips) only result in minor variations in outcomes. We run our simulations on the CPU on both machines and on the GPU on the second machine. We summarize our simulation results in Table 6. We focus on 1) a naive implementation which models what might be an expected “first-attempt” by an attacker, using the Python web3py library, and 2) the best performing implementation: coincurve + pycryptodome + multiprocessing.
The naive software implementation produces in the order of $10^{3}$$10^{4}$ addresses per second. The more optimized software implementation produces an order of magnitude improvement ($10^{4}$$10^{5}$). The Vanity-ETH implementation is much slower than the standalone clients. The CUDA implementation from Profanity2 is three orders of magnitude faster than our best CPU implementation with roughly half a billion APS. The Profanity2 developers self-report about a billion APS on similarly priced consumer hardware [^18] https://github.com/1inch/profanity2?tab=readme-ov-file#benchmarks---current-version.
@@
Based on the distribution of the matching digits $d$ for each attack group, we next attempt to infer the attacker’s hardware capability. We define two units: CPU-day and GPU-day, the number of addresses attackers can generate in one day per machine with a CPU and a GPU, respectively. From the above benchmarks, 1 CPU-day is $3.98\times 10^{10}$ addresses and and 1 GPU-day is $4.46\times 10^{13}$, respectively. In the process of producing a $d$-digit match, the attacker can generate many addresses matching $d^{\prime}<d$ digits in $\mathcal{R}$. Therefore, estimating the amount of computation needed to generate the maximum match $d$ for each group suffices. As shown in Figure 8, Group 1 produces a maximum 20-digit match while other groups (Groups 2, 3, and 4) achieve 14-digit matches. We calculate the number of expected trials for $d=20$ and, based on Table 4, $r=10^{6}$. We estimate that Group 1 have used $3.0\times 10^{7}$ CPU-days or 27,093 GPU-days. Group 1 would have needed to run over 41,000 CPUs around the clock over our whole two years of measurements to generate their lookalike addresses. We thus suspect they perform a substantial amount of their computation via GPUs. Other groups ($d=14$) only used 1.81 CPU-days, or $1.6\times 10^{-3}$ GPU-days.
To get a rough idea of the economic cost of computation, we use the AWS pricing model for similar GPUs and CPUs [^19] GPUs: https://aws.amazon.com/ec2/instance-types/g4/ and CPUs: https://aws.amazon.com/ec2/pricing/on-demand/. For GPU, an hourly rate (yearly plan) for g4dn.16xlarge (NVIDIA T4 tensor) is \$2.612 per hour; \$62.69 for one GPU-day. For CPU, an hourly rate (on-demand) for r6i.4xlarge is \$1.008; \$24.19 for one CPU-day. Using these numbers, we estimate that Group 1 would have paid (at most) 1.7M USD to generate a 20-digit match, which is below its 4M USD profit described in Table 5. Furthermore, these numbers are likely higher than the costs incurred by self-hosting machines. Notably, we do not consider specialized hardware for address generation (e.g., ASICs), which would further reduce the generation time/cost for Group 1. For other groups with $d=14$, CPU costs ($1.81\times 24.19\approx 43$ USD) are dominated by transaction costs.